how to prove cot x = 1/tan x by using trigonometric ratio?
Answer by Theo(13342)

let a = opposite side of angle x.
let b = adjacent side of angle x.

tanx = opposite side / adjacent side = a/b

cotx = adjacent side / opposite side = b/a

if you take 1 and divide it by tanx, you get 1 / tanx = 1 / (a/b)

since 1 / (a/b) is equal to 1 * (b/a), you get 1 / tanx = b/a

since b/a is equal to cotx, you get 1/tanx = cotx.

that's your proof.